### 02 Feb 2024 v1

# LLMs Can’t Plan, But Can Help Planning in LLM-Modulo Frameworks

Abstract
There is considerable confusion about the role of Large Language Models (LLMs) in planning and reasoning tasks. On one side are over-optimistic claims that LLMs can indeed do these tasks with just the right prompting or self-verification strategies. On the other side are perhaps over-pessimistic claims that all that LLMs are good for in planning/reasoning tasks are as mere translators of the problem specification from one syntactic format to another, and ship the problem off to external symbolic solvers. In this position paper, we take the view that both these extremes are misguided. We argue that auto-regressive LLMs cannot, by themselves, do planning or self-verification (which is after all a form of reasoning), and shed some light on the reasons for misunderstandings in the literature. We will also argue that LLMs should be viewed as universal approximate knowledge sources that have much more meaningful roles to play in planning/reasoning tasks beyond simple front-end/back-end format translators. We present a vision of LLM-Modulo Frameworks that combine the strengths of LLMs with external model-based verifiers in a tighter bi-directional interaction regime. We will show how the models driving the external verifiers themselves can be acquired with the help of LLMs. We will also argue that rather than simply pipelining LLMs and symbolic components, this LLM-Modulo Framework provides a better neuro-symbolic approach that offers tighter integration between LLMs and symbolic components, and allows extending the scope of model-based planning/reasoning regimes towards more flexible knowledge, problem and preference specifications.
1 Introduction
Large Language Models (LLMs), essentially n-gram models on steroids which have been trained on web-scale language corpora (or, effectively, our collective consciousness), have caught the imagination of the AI research community with linguistic behaviors that no one expected text completion systems to possess. Their seeming versatility has led many researchers to wonder whether they can also do well on planning and reasoning tasks typically associated with System 2 competency. On the face of it, this doesn’t seem to ring true, as both by training and operation, LLMs are best seen as giant pseudo System 1 Kahneman (2011) (see Figure 1). Even from a pure engineering perspective, a system that takes constant time to produce the next token cannot possibly be doing principled reasoning on its own. [^1] Think of asking an LLM an yes/no question–is this theorem logically entailed by this first-order logic knowledge-base. This is well-known to be a semi-decidable problem. Ask yourself if the LLM will take longer in answering the question. (If you are thinking Chain-of-thought prompts or training with step-by-step data, consider that you are essentially changing the nature of the original prompt/training). Not surprisingly, initial excitement based on anecdotal performance of LLMs on reasoning tasks Bubeck et al. (2023) has dissipated to some extent by the recent spate of studies questioning the robustness of such behaviors–be it planning Valmeekam et al. (2023c); Kambhampati (2023), simple arithmetic and logic Dziri et al. (2023), theory of mind abilities Ullman (2023); Verma et al. (2024), or general mathematical and abstract benchmarks McCoy et al. (2023); Gendron et al. (2023). Despite this, a steady stream of claims continue to be made in the literature about the planning and reasoning capabilities of LLMs.
In an ironic juxtaposition to this unwarranted optimism about the planning and reasoning abilities of LLMs, there is also unwarranted pessimism about the roles LLMs can play in planning/reasoning tasks. Several efforts (e.g. Liu et al. (2023); Pan et al. (2023); Xie et al. (2023)) advocate using LLMs only as glorified translators–for converting reasoning problems embedded in textual format to symbolic representations, which are then pawed off to external classical symbolic solvers (with all their attendant expressivity and search complexity challenges Doyle & Patil (1991)). [^2] In some circles, this unidirectional pipeline has come to be given the undeserved badge of neuro-symbolic architecture.
Figure 1: An informal account of viewing LLM as a giant external non-veridical memory that acts as a pseudo System 1
In truth, LLMs are a whole lot more than machine translators. They are a kind of approximate knowledge sources (albeit sans guarantees) trained on our collective consciousness. While it is unlikely that they will have System 2 competencies by themselves, they can nevertheless be valuable resources in solving System 2 tasks. The problem with Alchemy is not that Chemistry is useless, but that people wanted to delude themselves that Chemistry–a pretty amazing discipline on its own merits–can be Nuclear Physics if you prompt it just so. The confusions regarding LLM abilities, or should we say, LLM alchemy, seems to be not that much different–oscillating between ignoring what they are good at, and ascribing them abilities they don’t have.
The goal of this position paper is to introduce some clarity into this confusing state of affairs oscillating between over-optimism and over-pessimism. Simply put, we take the stance that LLMs are amazing giant external non-veridical memories that can serve as powerful cognitive orthotics for human or machine agents, if rightly used. The underlying n-gram nature makes them effortlessly intermix what would be considered disparate fields of study (not surprisingly, LLMs are seen to be very good at making/finding analogies!) The challenge is to leverage them without wrongly ascribing to them capabilities they don’t possess. The LLM-Modulo framework proposed in this position paper tackles this challenge.
For the sake of concreteness, we focus on planning tasks, especially as studied in the automated planning community Ghallab et al. (2004). The central position of the paper is that LLMs cannot plan themselves but can play a variety of constructive roles in solving planning tasks–especially as approximate knowledge sources and candidate plan generators in the so-called LLM-Modulo Frameworks in conjunction with external sound model-based verifiers.
We support this position by first reviewing literature that establishes that LLMs cannot be used as planners or plan verifiers themselves (Section 2). We also discuss why there are claims about planning/verification abilities in the first place, in the process hopefully clarifying some prevalent misunderstandings.
Second, we will propose a framework that allows us to leverage LLMs effectively in planning tasks, by combining them with external critics, verifiers and humans. We call this an LLM-Modulo Framework (a name loosely inspired by SAT Modulo Theories Nieuwenhuis & Oliveras (2006)); see Figure 3. LLMs play a spectrum of roles in this architecture, from guessing candidate plans, to translating those plans into syntactic forms that are more accessible to external critics, to helping end users flesh out incomplete specifications to helping expert users acquire domain models (that in turn drive model-based critics). All this leveraging of LLMs is done without ascribing to them any planning verification abilities–thus ensuring that the plans generated in this architecture can have formal correctness guarantees where possible.
One question regarding this LLM-Modulo architecture from a planning perspective is whether it is more han a gratuitous attempt to shoe-horn LLMs to solve planning problems, when there are more formal combinatorial planning systems available in multiple communities Ghallab et al. (2004). Compared to a planner that is guaranteed to be correct in a narrow set of domains, LLMs may likely be good at generating plausible (but not guaranteed to be correct) plan heuristics/suggestions in many more scenarios. Thus, unlike the traditional planning architectures studied in AI Ghallab et al. (2004), which put a priori constraints on the expressiveness of the problems that can be posed to the planner (to wit, the different expressiveness levels of the PDDL specification McDermott et al. (1998)), the LLM-Modulo architecture puts no such restrictions. In this sense, it is more representative of real-world planning problems such as those in NASA mission planning, where the different critics–human and automated–are at best able to give “no objection” certificates, clearing it from their perspective. (Indeed, both deep space network planning and mars rover task planning are done via a collective human blackboard. Johnston et al. (2014) Bresina et al. (2004).) Note that this is starkly different from just sending an unvetted plan out to execution (as would be the case if we have LLMs operate in autonomous mode to guess plans). Generalizing planning and reasoning frameworks this way is consistent with the Doyle & Patil’s call to the Knowledge Representation community of yore Doyle & Patil (1991).
2 Planning-centered Limitations of LLMs
In this section, we will first review the literature that calls into question the claims about planning and self-verification capabilities of LLMs. Subsequently, we will also provide some possible reasons for the claims to the contrary made in the literature.
2.1 LLMs cannot generate executable plans in autonomous mode
Despite the initial claims about the planning capabilities of LLMs Bairi et al. (2023); Yao et al. (2023b); Shinn et al. (2023); Huang et al. (2022); Hao et al. (2023) several recent studies independently confirm that LLMs are not actually able to generate executable plans when they are used in autonomous modes Valmeekam et al. (2023c); Liu et al. (2023); Silver et al. (2022). For example, in Valmeekam et al. (2023c, b), the authors evaluate LLMs ability to generate correct plans on a suite of planning problem instances based on the kinds of domains employed in the International Planning Competition IPC (1998). To eliminate the subjective aspect of analysis that forms the core part of many earlier efforts on evaluating the reasoning capabilities of LLMs, they automate the evaluation by leveraging models and tools from the automated planning community.
They show that the results in the autonomous mode are pretty bleak. On average, only about 12% of the plans that the best LLM (GPT-4) generates are actually executable without errors and reach their goals. They show that the choice of the specific LLM (they have tested the family of GPT LLMs including GPT-4 OpenAI (2023), GPT-3.5 OpenAI (2022), InstructGPT-3 Ouyang et al. (2022) and GPT-3 Brown et al. (2020)), as well as fine-tuning does not seem to have a major effect on this dismal performance. They also show that the performance deteriorates further if the names of the actions and objects in the domain are obfuscated–a change that doesn’t in any way affect the performance of the standard AI planners. This latter further suggests that LLMs are more likely doing approximate retrieval of plans than planning.
2.2 LLMs cannot verify plans and thus cannot improve by self-critiquing
There still exists considerable optimism that even if LLMs can’t generate correct solutions in one go, their accuracy might improve in an iterative prompting regime, where LLMs will be able to ”self-critique” their candidate solutions and refine them to the point of correctness Yao et al. (2023b, a); Shinn et al. (2023); Weng et al. (2023); Huang et al. (2022). This belief seems to rest largely on the assumption that verification of correctness should be easier than generation for many reasoning problems–a rather classical argument from computational complexity. There are grounds to be skeptical of this assumption as the complexity of the reasoning task should be irrelevant to LLM performance if what they are doing is approximate retrieval. In general, unless LLMs are trained not just on “correct data,” but also “corrections data,” there is no a priori reason to believe that their critiques would even be approximately relevant, let alone actually correct.
Two recent studies–one on plan verification Valmeekam et al. (2023a) and the other on CSP verification Stechly et al. (2023) seem to throw cold water on this optimism. In Stechly et al. (2023), the authors systematically investigate the effectiveness of iterative prompting in the context of Graph Coloring, a canonical NP-complete reasoning problem. Their methodology involves a principled empirical study of the performance of GPT4 on two tasks: solving a large suite of random graph coloring instances and, separately, verifying the correctness of the candidate colorings–both in direct (i.e., return the first solution generated by the LLM) and iterative modes. In iterative modes, they experiment both with an LLM critiquing LLM-produced solutions and an external, guaranteed correct reasoner verifying solutions. In both cases, they analyze whether the content of criticisms actually affects bottom-line performance.
Their results indicate that in direct mode, LLMs are, perhaps not surprisingly, pretty bad at solving graph coloring instances. More interestingly, they are no better at verifying solutions. In iterative modes, given the inability of LLMs to verify solutions, it should come as no surprise that their experiments also show that the strategy of LLMs self-critiquing their solutions does not improve over the baseline. It is apparently worse because the system can’t recognize a correct coloring and thus merrily passes over fortuitously correct colorings it has generated, ending up with a wrong one! Similar results have also been reported for planning problems in Valmeekam et al. (2023c).
One important corollary of the fact that LLMs cannot self-critique their plans is that they can’t also self-improve by generating synthetic data by generating plans themselves, critiquing the plans by themselves to improve them, and then use those to fine-tune themselves, as has been claimed in the literature Wang et al. (2022) [^3] Contrary to the claim of “self-improvement”, works like Wang et al. (2022) actually heavily depend on external knowledge (crafted seed examples) and critics (filtering step). Huang et al. (2023b); see Section 3.3.
2.3 Analyzing Claims to the Contrary in the Literature
Given that LLMs can neither guarantee correct generation nor correct verification of plans, as discussed in the previous sections, one obvious question is why the literature is replete with claims contrary to this Bairi et al. (2023); Yao et al. (2023b); Shinn et al. (2023); Yao et al. (2023a); Weng et al. (2023); Huang et al. (2022).
Claims about Planning: To analyze the planning claims, we need to first understand that solving planning tasks requires (a) having the necessary planning domain knowledge–the actions and their preconditions, effects; the standard hierarchical recipes (e.g. task reduction schemas in HTN planning), past cases/plans, etc., and (b) being able to assemble this planning knowledge into an executable plan that takes care of any subgoal/resource interactions. The first can be called the knowledge acquisition and the second reasoning/planning part. Many of the papers claiming planning abilities of LLMs, on closer examination, wind up confusing general planning knowledge extracted from the LLMs for executable plans. When all we are looking for are abstract plans, such as “wedding plans,” with no intention of actually executing the said plans, it is easy to confuse them for complete executable plans. Indeed, our close examination of several works claiming planning capabilities for LLMs suggests that they either work in domains/tasks where subgoal interactions can be safely ignored Yao et al. (2023b); Shinn et al. (2023) [^4] Although domains like AlfWorld Shridhar et al. (2021) do have sub-goal interactions for successful task completion, Yao et al. (2023b) and Shinn et al. (2023) ignore these interactions relying on the ergodic nature of the domain when prompting LLMs for generating plans. –either because they are just working on a single goal, or because the world is forgiving and ergodic; or delegate the interaction resolution (reasoning) to the humans in the loop (who, through repeated prompting, have to “correct” the plan). Sometimes, in common sense domains, or with enough fine-tuning, the “assembling” part may also be obviated by having seen a case that pretty much corresponds to the problem that needs to be solved. Not surprisingly, the work by Valmeekam et al. (2023c) shows that if the action interactions are removed by relaxing the world models, then the ability of LLMs to guess executable plans improves. Without these assumptions or mitigations, the plans that come out of LLMs may look reasonable to the lay user, and yet lead to execution time interactions and errors. [^5] These issues are illustrated in part by a recent news story Kugel & Hiltner (2023) about the proliferation of travel planning books, mostly auto-extracted from LLMs, and the ensuing disappointment of the unsuspecting end users who buy them mistaking them for usable plans!
The fact that LLMs are often good at extracting planning knowledge can indeed be gainfully leveraged. As shown in recent works Guan et al. (2023), LLMs can be a rich source of approximate models of world/domain dynamics and user preferences, as long as the humans (and any specialized critics) in the loop verify and refine those models, and give them over to model-based solvers. This way of using LLMs has the advantage that the humans need only be present when the dynamics/preference model is being teased out and refined, and the actual planning after that can be left to sounder planning frameworks with correctness guarantees, such as LLM-Modulo.
Figure 2: Viewing LLMs as an approximate knowledge source trained over civilizational knowledge
Such an overall approach has striking similarities to knowledge-based AI systems of yore, with LLMs effectively replacing the “knowledge engineer” (see Figure 2). Given the rather quixotic and dogmatic shift of AI away from approaches that accept domain knowledge from human experts that some writers termed “Polanyi’s Revenge” (c.f. Kambhampati (2021)), this new trend of using LLMs as knowledge sources can be viewed as a form of avenging Polanyi’s revenge! Indeed, LLMs make it easy to get problem-specific knowledge as long as we are willing to relax the correctness requirements of that knowledge. In contrast to the old knowledge engineering approaches, LLMs offer this without making it look like we are inconveniencing any specific human (we are, instead, just leveraging everything humans told each other!). So the million dollar question for reasoning tasks is: ”how would you do robust planning if you have some doddering know-it-all ready to give you any kind of knowledge?” The LLM-Modulo Framework is a principled method for tackling this challenge.
Figure 3: The proposed LLM-Modulo framework where LLMs act as idea generators and various external critiques that specialize in different aspects, critique the candidate plan.
Claims about Self-Verification: Coming to the claims about LLM’s self-verification abilities, a closer look at the literature Yao et al. (2023a); Huang et al. (2023a) shows that those claims are either (i) made in the context of tacit knowledge tasks for which there is little possibility of a verifier (e.g. essay writing)–making it hard to evaluate whether LLM’s critiquing actually helped [^6] Paradoxically, the fact that it is infeasible to write sound verifiers for tacit knowledge tasks also makes it possible for everyone to be a critic. Think of R2 saying the paper could be made ”less dense” or the Peloton instructor critiquing Christopher Nolan film. or (ii) the external verification is carried out either by simulators Wang et al. (2023); Yao et al. (2023b) or simple calls to the underlying operating system (as is the case, for example, for the 24 puzzle in Yao et al. (2023a)).
In a related vein, there is the recent Tree of Thoughts (ToT) paper Yao et al. (2023a), which has been pitched as a way to convert LLMs into some type of systematic search with self-verification. A closer look at the work however shows that ToT simply iteratively back-prompts the LLM until it comes up with a solution that is acceptable to an external verifier. Specifically, ToT employs a problem-specific prompt priming method. The “tree” in ToT is essentially a way to generate diverse priming prompts (that the authors set up in a problem specific way). In other words, despite the use of terminology of problem-solving agents Russell & Norvig (2010)–search tree, expansion etc., there is no deeper connection to search-based agents. [^7] Our preliminary experiments also show that at least in 24 puzzle, a simple iterative prompting, even without a systematic prompt diversification, is quite competitive with the ToT framework.
The guarantees–if any–are coming in terms of soundness of the external verifier. The one clear reasoning problem used in the ToT paper is the 24 puzzle–for which the external verifier can be easily implemented in terms of arithmetic operations (thankfully not done by the numerically challenged LLM!). Here, our experiments show that LLM’s own criticisms are often quite off the mark. [^8] Note that we can do this check easily because of the formal specification of correctness. For the “improving writing task” also used in ToT, there are no formal quality metrics and so it is hard to say anything concrete about the critiques of the LLM. Because the 24 puzzle’s solutions can be verified by simple arithmetic operations, readers don’t quite realize that the framework relies on an external verifier. In general though, the verifier may be more complex and can involve substantial work (you can substitute a simulator for the verifier–but someone has to write that simulator too!)
In general planning problems, one way to provide an external verifier is to (a) write a domain model (e.g. in PDDL) and (b) feed it to an off-the-shelf model-based verifier like VAL (c.f. Guan et al. (2023)
3 LLM-Modulo Framework for Robust Planning
While Section 2 questions the claims that LLMs are capable of planning/reasoning by themselves, it is certainly not meant to imply that LLMs don’t have any constructive roles to play in solving planning/reasoning tasks. On the contrary, as discussed in the Introduction, their uncanny ability to generate ideas/potential candidate solutions–albeit with no guarantees about those guesses–can be valuable in the generate-test-critique setups in conjunction with either model-based verifiers or expert humans in the loop. Accordingly, we propose a general “LLM-Modulo” framework [^9] The name LLM-Modulo is inspired by the SAT-Modulo theories Nieuwenhuis & Oliveras (2006). While we believe that versions of such an architecture can be of use in a wide variety of planning or reasoning tasks, for the sake of concreteness, we will focus on planning tasks, especially of the type studied in the automated planning community Ghallab et al. (2004).
Figure 3 gives a schematic of the LLM-Modulo Framework, as we envision it. As can be seen readily, the underlying architecture is a Generate-Test-Critique loop, with the LLM generating candidate plans and a bank of critics critiquing the candidate. The loop starts with the LLM getting the problem specification and generating its first plan candidate. [^10] Although we focus on planning from scratch, it is easy to accommodate replanning scenarios, where the loop starts with an externally supplied candidate plan. Note that the plans an LLM helps generate in this architecture have soundness guarantees because of the external sound critics. This means that plans coming out of such an architecture will constitute a better corpus of synthetic data for any fine tuning phase carried out to improve/customize the LLM’s generation capability.
Design Considerations: Before going into the details about the framework and its various modules, it is worth noting some design decisions underlying the proposed architecture. We start by noting that the LLM-Modulo architecture is a “Generate-Test” one that involves LLMs interacting with the external critics rather than solvers. This is a deliberate decision–as this allows the LLM to guess/generate candidates to satisfy the critics, as against dealing with the expressiveness and search complexity issues of the solvers. Secondly, the framework explicitly recognizes that the LLMs can generate approximate ideas not just about plan candidates, but domain models, problem reduction strategies, and refinements to the problem specification. The framework also recognizes that LLMs are good at format/syntax changes. Accordingly, the framework leverages all these abilities of LLMs, letting them play multiple roles in planning. Finally, the architecture carefully circumscribes the human’s role–domain experts interact with the LLM to tease out the models used by (some of) the critics, while end users take part in refinining any incomplete problem specification in concert with the LLM. A notable, and deliberate, absence is human’s involvement in the inner loop of planning–e.g. with iterative prompting. In addition to posing an infeasible burden on the human’s time for complex planning problems, such iterative prompting strategies are notorious for their Clever Hans effect cle .
3.1 Critics
In the LLM-Modulo framework, critics can evaluate LLM-generated candidates for a planning/reasoning problem over both hard and soft constraints. Hard constraints refer to correctness verification which can include causal correctness, timeline correctness, resource constraint correctness, etc. For PDDL planning problems, the hard critic can be based on VAL Howey et al. (2004), that works off of a model (which itself can be acquired with the help of the LLM Guan et al. (2023). On the other hand, soft constraints can include more abstract notions of correctness such as style, explicability, preference conformance, etc. As discussed in Section 2.3, while LLMs cannot take on the role of hard critics, they can help simulate some aspects of the role of soft critics. So our framework does allow for style critics be possibly based on LLMs (e.g Verma et al. (2024)). We reiterate that the soundness of the LLM-modulo framework is inherited from the soundness of the critics.
Critics can range from those that check for the hard constraints of correctness–causal correctness, timeline correctness, resource constraint correctness etc., to those that check for soft constraints such as style, explicability, preference conformance etc. As discussed in Section 2.3, while LLMs cannot take on the role of hard critics, they can help simulate some soft critics (e.g. Verma et al. (2024)). For PDDL planning problems, the hard critic can be based on VAL Howey et al. (2004), which works off of a model (which itself can be acquired with the help of the LLM Guan et al. (2023).
The bank of critics–hard (model-based) as well as soft (possibly LLM-based) evaluate the current plan candidate to evaluate its fitness/acceptability. If at least all the hard critics sign off on the current candidate, then that is considered a valid solution to be returned to the end-user or the executor. When a critic finds the current plan candidate to be unsatisfactory, it can provide varying levels of feedback, ranging from “No, try again” to “No, try again, here is one thing wrong with the current plan” to “No, try again, here are all the things wrong with the current plan. These critiques are all pooled at the Backprompt Controller (see Section 3.2)
3.1.1 LLMs as Reformulators
One interesting challenge is that many of the symbolic model-based verifiers tend to be operating on specialized formal representations. Given a central candidate plan (e.g. a mission plan), these critics need translations of that candidate into their representations. This is the role of the reformulator module attached to individual critics. These reformulator modules can be supported to large extent by LLMs, given that one thing LLMs are very good at is format change across different syntactic representations, Olmo et al. (2021). Indeed, as discussed in the Introduction, some approaches to combine LLMs with external symbolic solvers just use LLMs as reformulators for these solvers Liu et al. (2023); Pan et al. (2023). Our discussion of LLM-Modulo framework should make it clear that syntax reformulation alone is a severely limited role for LLMs!
3.2 Backprompt (Meta) Controller
The critiques from the various critics are pooled together by the Meta (Backprompt) Controller, which passes a processed version of them to the LLM as the next iterative prompt to elicit the next guess. This is especially required in the presence of a mix of soft and hard critics, where the Meta Controller can assume the responsibility of compiling the critiques into a consistent feedback to process.
The processing steps taken in the controller can range from simple round-robin selection of prompts to generating a summarized prompt (with LLM help) to employ a prompt diversification strategy to elicit the next candidate from a different part of the implicit search space (akin effectively to the strategy used in systems such as Tree of Thoughts prompting Yao et al. (2023a), as discussed in 2.3).
3.3 Fine Tuning & Synthetic Data
Once the LLM-Modulo framework “solves” a planning instance, the solution can then be added to a synthetic data corpus (step 6 in Figure 3), which is intermittently used to fine tune the LLM (step 7), so its future plan candidate guesses improve.
Such fine tuning on task-specific data has been a popular way to get LLMs to improve their performance on reasoning/planning tasks. For example, fine tune the LLM on blocks world planning problem-solution pairs to improve their performance in guessing solutions for blocks world instances Pallagani et al. (2023). While fine tuning still doesn’t guarantee correctness of the generated solutions, it might improve the chances that LLM guesses candidates that are closer to being vetted by the bank of critics.
One important question is where this additional data for fine tuning comes from. A tempting idea is to have the LLM itself generate this additional data, and improve it by self-critiquing/verification, before fine-tuning itself on the data. This EM-like approach unfortunately will not work given that LLMs can’t verify their own solutions (see Section 2.2). In the past, this meant that the only reliable way to generate synthetic data is to use external plan generators–for example, use a classical planner like FF to solve a host of blocks world instances and then use those solutions to fine-tune the LLM. The LLM-Modulo framework, in contrast, provides an alternative way of using an LLM-based framework to generate synthetic data that is guaranteed correct.
3.4 Specification Refinement & Model Acquisition (Semi-automated)
As mentioned earlier, we avoid having humans involved in iteratively prompting LLMs–as this can be an infeasibly time-consuming activity for them. Instead, we let automated verifiers, either model-based or LLM-supported, to manage the plan critiquing process. The framework does depend on humans for “once per domain” and “once per problem” critiques. In the former category, human domain experts can play a role in acquiring the domain model with the help of the LLM. Examples of such interaction include teasing out PDDL planning models from the LLMs with the help of human expert curation (top left in Figure 3). The idea here is that the traditional domain model acquisition task (e.g. sim (2001)) is significantly made easier by having the LLMs help with ideas regarding various pieces of the domain model (e.g., actions, their preconditions and effects) and letting humans sign off/critique the resulting model. Once the model is acquired this way, it can be used by correctness verifiers such as VAL Howey et al. (2004); Guan et al. (2023). Often the planning problems in real world situations are specified incompletely, leaving it to the human commonsense to refine the specification. This brings up a second role for humans–this time end users (bottom left in Figure 3–in collaboratively refining the specification with the help of LLMs (similar to the way done in Xie et al. (2023); Liu et al. (2023)).
3.5 Summary of Roles of LLMs in the LLM-Modulo Framework
It is worth summarizing the multiple roles the LLM plays in the LLM-Modulo architecture. The most prominent, of course, is its role in “guessing” the candidate plans (step 2 in Figure 3) in response to problem specification and iterative back prompting from the bank of critics (Step 5). Second the LLM plays a role in converting the guessed plan candidate into specialized representations used by the various critics (e.g., the time-line view, the causal link view etc.). This role leverages the fact that LLMs are very good at format conversion (c.f. Olmo et al. (2021)) Third, the LLM plays a role in helping the end user flesh out the incomplete problem specification to begin with (Step 1 in Figure 3). Finally, the LLM plays a role in helping the domain expert tease out and refine the domain models used by the various model-based critics Guan et al. (2023); Kwon et al. (2022).
4 Related Work
We will note that while the LLM-Modulo framework is being proposed in general form here for the first time, there are certainly works in leveraging LLMs in planning and reasoning tasks that are in line with the spirit of the LLM-Modulo framework. For exmaple, both Valmeekam et al. (2023c) and Stechly et al. (2023) describe and evaluate a backprompting interaction between an LLM and an external verifier. Work on FunSearch Romera-Paredes et al. (2023) depends on a generate-test loop between a specially fine-tuned LLM that guesses solutions, and an external symbolic evaluator that critiques them. The authors note how the external verifier is critical for avoiding falling prey to hallucinations (i.e., approximate solution candidates that have flaws). AlphaGeometry Trinh et al. (2024) too depends on the Generate-Test-Critique interaction between a fine-tuned LLM and a symbolic evaluator. Both these systems fine-tune pre-trained LLMs with task specific synthetic data–the correctness of which is vetted with external simulators (as we discuss in 3.3).
While we focused on PDDL planning tasks for the sake of concreteness, we believe that the essence of LLM-Modulo framework is equally applicable to other scenarios involving planning and reasoning–such as Reinforcement Learning with Simulators. Such RL systems rely on rewards as feedback to train a policy. Analogous to the plan evaluation and critiques performed by the respective critics in the LLM-Modulo framework, environment simulators provide the required feedback. This has particularly been useful but largely overlooked in cases where LLMs are used as an actor to generate an admissible plan by interacting with a simulator, for example in the case of AlfWorld Yao et al. (2023b); Shinn et al. (2023) and Minecraft Wang et al. (2023). As mentioned in Section 3, similar to extracting a domain model such as in the case of PDDL planning Guan et al. (2023), designing a reward model for the plan generation - feedback cycle is yet another potential use case that has been recently looked at for text-based Kwon et al. (2022); Hao et al. (2023) and robot manipulation Ma et al. (2023) domains.
Interestingly, the fact that LLM’s can help come up with approximate quasi-symbolic transition models, reward models and models of high level actions has made a bigger splash in RL. This is because for far too long, researchers there have tried to spurn any high level models (lest that would involve depending on humans) and focused on learning to act from sensory information, under the name of “deep reinforcement learning.” Given the horrendous sample complexity of the DRL methods even in reaching a single subgoal, and the well known fact that even approximate symbolic models can help drastically improve the performance (c.f. Guan et al. (2022)), coupled with the fact that LLM’s are only too glad to dream up approximate models and goal recipes, there has been a performance revolution of sorts there Yao et al. (2023b); Liang et al. (2023); Wang et al. (2023). If we look beyond the improvements in these lower level goal seeking behaviors–especially in the presence of ergodic simulators, the RL approaches dependent on LLMs will encounter the same issues regarding subgoal interactions that our discussion of PDDL planning problems brought into focus. The LLM-Modulo inspired frameworks will thus, we believe, be equally relevant there. Indeed, SayCan Ahn et al. (2022) the earliest use of LLMs in generating policies in an RL-with-Simulator scenario, explicitly filters the action choices suggested by the LLM with the help of simulator.
While we focused on text based LLMs (such as GPT4), recently there have also been impressive development in multi-modal LLMs (e.g. GPT4V). While multi-modality is a great addition that increases the coverage of their System 1 imagination (Figure 1), it is not clear that this gives them System 2 competence.  [^11] If you know how to complete sentences, and now learned to complete dance moves, does your ability to reason/plan magically improve?
5 Conclusion
This position paper is a modest attempt to combat both over-optimism and over-pessimism about the role of LLMs in planning and reasoning tasks. Our position is that LLMs cannot plan themselves but can play a variety of constructive roles in solving planning tasks–especially as approximate knowledge sources and candidate plan generators in the so-called LLM-Modulo Frameworks in conjunction with external sound model-based verifiers. In support of this position, we summarized the literature questioning the claims about the planning and self-verification capabilities of LLMs by themselves. We also discussed how conflating approximate knowledge acquisition and generating executable plans of action are behind many of the claims about planning and verification abilities of LLMs. We then shared LLM-Modulo framework, our vision for a productive way to integrate the impressive idea generation/approximate knowledge provision capabilities of LLMs with external verifiers with correctness guarantees for robust planning. We discussed how planning in LLM-Modulo framework avoids inheriting the expressiveness and search-complexity issues of traditional symbolic planners, while retaining their soundness guarantees. As we discussed, LLM-Modulo frameworks are consistent with some of the most high-profile success stories of “neuro-symbolic” architectures, including AlphaGeometry and FunSearch.